OpenStudy (anonymous):

What are the solutions of the system? y=x^2+3x-4 y=2x+2

OpenStudy (anonymous):

answers: (–3, 6) and (2, –4) (–3, –4) and (2, 6) (–3, –4) and (–2, –2) no solution

OpenStudy (earthcitizen):

\[x ^{2}+3x-4=2x+2 => x ^{2}+x-6=(x-2) (x+3)=0\] x=2 or -3 at x = 2, y=0, x=-3, y=0

OpenStudy (shayaan_mustafa):

@EarthCitizen Help him please. I am a little bit busy

OpenStudy (shayaan_mustafa):

Thanks

OpenStudy (earthcitizen):

no solution

OpenStudy (earthcitizen):

no p @Shayaan_Mustafa

OpenStudy (phi):

to solve y=x^2+3x-4 y=2x+2 if both equal y then they equal each other x^2+3x-4= 2x+2 as posted above, this simplifies (add -2x-2 to both sides) x^2 +x - 6=0 this factors to (x+3)(x-2)=0 of course this gives two roots: x=2 and x=-3 to find y, use either of the original equations. Best to use the simple one: y= 2x+2 if x= 2, y= 2*2+2= 6 so (2,6) is one of the solutions. I'll leave it to you to find the 2nd solution (when x= -3)

OpenStudy (earthcitizen):

kk, my bad.

OpenStudy (anonymous):

so what is the answer??

OpenStudy (earthcitizen):

@phi when either x is used to find y, does it give the same values for y ?

OpenStudy (earthcitizen):

(2,6) and (-3,-4)

OpenStudy (phi):

@EarthCitizen Yes. that is the whole idea.